OpenStudy (anonymous):

Find f(x) and g(x) so that the function can be described as y = f(g(x)). y =

OpenStudy (anonymous):

f(g(x))

OpenStudy (anonymous):

sorry Y=

OpenStudy (anonymous):

F(x) = 9/x g(x) = sqrt (5x+5)